# Likelihood-based Sensor Calibration for Expert-Supported Distributed Learning Algorithms in IoT Systems

### v1

@@ III-B Multivariate Regression Approach
Algorithm 3 Hybrid solution: Implementation of a combination of the simple MLE and Gleser et al.
@@
[…]
2:  $[\mathbf{V},\mathbf{D}]=\mathrm{eigs}(\mathbf{X}^{T}\mathbf{X}+\mathbf{Y}^{T}\mathbf{Y})$
3:  $\boldsymbol{\Theta}^{T}=\mathbf{V}\mathbf{V}^{T}\mathbf{X}$
4:  $\mathbf{B}=\mathbf{Y}\mathbf{X}^{T}(\mathbf{X}\mathbf{X}^{T})^{-1}$
5:  $\tilde{\mathbf{A}}=\mathbf{B}(1:p-1,1:p-1)$
6:  $\tilde{\mathbf{b}}=\mathbf{b}(1:p-1,p)$
7:  $\boldsymbol{\Theta}_{E}=\boldsymbol{\Theta}(1:p-1,:)$;
TABLE II: Normalized transformation error $\bar{e}_{y}$ linearly scaled ($min=0.00072817$ resulting 0, $max=0.0028$ resulting 1) using real data. Averaged comparison for each sensor dataset with the transformed data from all sensors for the implementation of Gleser et al., Alg. 1, Alg. 2, Alg. 3, and the feature-wise normalized dataset.
@@
IV Results
We evaluate the presented algorithms and investigate the difference between the implementation of Gleser et al. and our denoised augmented improvement using Monte Carlo simulations. Then, to test the applicability in practice, we feed the algorithms with data from an artificial nose application.
IV-A Monte Carlo Simulations
We developed Monte Carlo simulations to evaluate the algorithms. In $l=1000$ simulation runs $n=1000$ data vectors $\boldsymbol{\theta}_{i}$ were generated for standard deviations $\sigma$ from $1$ to $15$ using a data generation tool from [23]. The data vectors have a dimension of $q=2$. We used a fixed transformation given by:
where $\tilde{\boldsymbol{\theta}}=[\tilde{\boldsymbol{\theta}}_{1},...,\tilde{\boldsymbol{\theta}}_{l}]$.
The solution of Gleser et al. also estimates $\boldsymbol{\Theta}$, therefore, we also calculate the error defined in (17):
Finally, we calculate the arithmetic mean values of the errors $e_{x}$ and $e_{y}$ over the $l=1000$ simulation runs. The arithmetic mean values are denoted with $\bar{e}_{x}$ and $\bar{e}_{y}$. Figure 6 shows the mean estimation errors $\bar{e}_{x}$ and $\bar{e}_{y}$ of the algorithms.
Since Algorithm 2 only estimates the parameters of the transformation between the noisy measured data from system 1 to system 2, it performs worse than Gleser et al. taking the origins into account. In this case, the error $\bar{e}_{x}$ is only the noise of system 1. The hybrid version in Alg. 3 benefits from Gleser et al. and delivers same quality results.
Analyzing error $\bar{e}_{y}$ in table I shows the improvement of the modification. Alg. 1 improves the results of Gleser et al. with rising standard deviation $\sigma$ and Alg. 3 delivers another slight improvement. Even the simple least squares solution of Alg. 2 can perform better than Gleser et al. looking at the standard deviations greater than 7.
IV-B Gleser augmentation eigenvector denoising
Figure 5 shows the results of the Monte Carlo simulation for the eigenvector denoising extension of Alg. 1 in comparison with the Gleser et al. standard implementation. One can see the improvement in the estimation of $\bar{e}_{y}$ while error $\bar{e}_{x}$ remains the same.
Table I gives an overview of the percentual mean errors $\bar{e}_{x}$ and $\bar{e}_{y}$ for the implementation of Gleser  et al., Alg. 1, Alg. 2, and Alg. 3 for the selected values of the standard deviation $\sigma$.
@@ IV-B Gleser augmentation eigenvector denoising
IV-C Real Data example 8 sensor board
For the collection of real test data we measured effects of humidity, temperature and disturbances in the air. In addition, we heat the samples during data recording and clean the sensors at regular intervals. The Bosch BME688 development kit board, shown in figure 3, is equipped with eight identical BME688 metal oxide sensors that are used for data recording.
For the experiment, we prepared two samples [26], water and apple juice, each with a volume of 100 ml. We present only the apple juice investigations for clarity. The results with water are similar and are available in the gitlab repository. Before each measurement of the sample, the sensor is cleaned in air. The measurement period per glass is 20 minutes. The temperature of the sample glass is increased by 1 °C every 2 minutes during the measurement. The total temperature increases from 25 °C at the beginning of the measurement, to 35 °C at the end of the measurement period. The Bosch BME688 is operated with a continuously repeating temperature cycle of 10 heating temperature steps, also called parallel mode. In this example, the temperature cycle is set to 5 steps at 200 °C and 5 steps at a sensor temperature of 400 °C, each at an interval of 1260 milliseconds. For later evaluation of the measurement data, the arithmetic mean is calculated from the 5 raw values at 200 °C and 400 °C.
As already mentioned in section I, figure 2 shows the results of the estimated transformation for an normalized example data set in the use-case of an artificial nose measuring apple juice. The black dots show the position of the data points in system $1$, blue dots represent the measured data in system 2 and the red dots are the estimation of system 2 based on the calculated transformation of Proposition 1. It can be seen that Proposition 1 allows a good estimation of the transformation in this case.
Figure 7 shows the results of the transformation error $e_{y}$ for the normalized data set of the multi-sensor board. We calculated the error ${e}_{y}$ for the $64$ combinations of sensors denoting $j$ as the source sensor and $k$ as the target sensor, $j,k\in\{1,...,8\}$, by (18)
where the $\tilde{\mathbf{A}}_{j,k}$ and $\tilde{\mathbf{b}}_{j,k}$ are representing the transformation parameters from source sensor $j$ to target sensor $k$. The mean error for the $n=90$ data samples is calculated by (19)
$\bar{e}_{y}=\frac{1}{n}\sum_{i=1}^{n}\lVert(\tilde{\mathbf{A}}_{j,k}{\mathbf{x}_{j}}+\tilde{\mathbf{b}}_{j,k})-\mathbf{y}_{k}\rVert_{2}.$ (19)
Since we can use the measured values $\mathbf{y}_{k}$, we do not need to calculate them as in eq. 16 by $(\mathbf{A}\boldsymbol{\theta}_{i}+\mathbf{b})$.
Table II shows the normalized mean error $\bar{e}_{y}$ for each source sensor, with the minimum value $0.00072817$ resulting 0, and maximum value $0.0028$ resulting 1. Since we focus on calculating the transformation from $\mathbf{X}$ to $\mathbf{Y}$, and do not try to estimate the origins of the data, the performance of Gleser et al. is slightly less accurate than the least squares, and the hybrid version delivers same results as Proposition 1. Our proposed Alg. 2, with reduced computational steps and therefore lower energy consumption, gives the best results except source sensor 6. However, this could also be due to inaccuracies in the measurement procedure and needs further investigation. A simple normalization of the data in our experiment leads to at least $50~{}\%$ worse results than our proposed method.
@@ V Discussion and Outlook
In the article, we have shown that the procedures of estimating the 2-dimensional AT can be used for an expert-assisted learning between two systems. The simple direct estimation based on the stationary points of the parameters of the transformation can be combined with the solution for the estimation of the data from Gleser et al.. This new hybrid solution provides the best estimates for both, simulations and real data. For the presented real data evaluation, we achieve best results with the simple Prop. 1, but we are looking forward to investigate whether we can use the approach of Gleser et al. to improve the measurement values of our low-cost sensors in an extended setup. For the real data example, our simple least squares approach already achieves unique solution due to the non singularity in $\boldsymbol{\Theta}\boldsymbol{\Theta}^{T}$, and the best accuracy with low computation, which is very promising for further developments in distributed learning. The results still need to be generalized for the m-dimensional case. Finally, in further work we will explore if only a part of the measured data is sufficient to give a good estimate of the transformation. Based on this, iterative methods can be developed in the future to estimate the transformation between different systems. As to our knowledge, there is no similar approach, although it seems to be a simple way to match distributed systems and their events or anomalies. Figure 8 shows the idea of transforming similar sensor data to create similar models using the expert-supported transformation estimation (see figure 1) between the two data spaces. We plan to implement these concepts as the basis for a more efficient method of knowledge sharing between distributed nodes.
@@ VI Conclusion
We propose an improved solution of Gleser et al. by augmentation of the matrix, denoising using the eigenvectors and the extension to an AT. We have shown that both solutions are connected by an eigenvalue decomposition. We further showed that a combination of both solutions provides better estimates for computing the parameters of the AT by both, simulation, and a real experiment. Finally we presented a novel concept for distributed learning using transformation with the support of expert knowledge.
All 3 algorithms offer a good approach to estimate the AT between similar systems. Further investigations should include special methods to fairly compare Gleser et al. with our hybrid proposition. The development of a simple but powerful expert-based distributed learning algorithm using ATs appears promising but requires further research in multidimensionality, and an iterative approach. Investigating more powerful transformation for stronger IoT nodes should make sense as well.
Acknowledgment
This project was partly funded by the German Federal Ministry of Food and Agriculture (BMEL) projects ”KI-Pilot” under Grant 2820KI001, and ”PINOT” under Grant 28DK107G20, and the German Federal Ministry for Economic Affairs and Climate Action (BMWK) project ”EASY” under Grant 01MD22002D, and the German Federal Ministry for the Environment, Nature Conservation, Nuclear Safety, and Consumer Protection (BMUV) project ”KIRA” under Grant 67KI32013B, and the Ministry for Science and Health of Rhineland-Palatinate (MWG) as part of the research training group (Forschungskolleg) ”AI-CPPS” https://aicpps.uni-trier.de.
@@ Acknowledgment
$\nabla_{\boldsymbol{\Theta}}\mathrm{tr}[(\mathbf{X}+\boldsymbol{\Theta})(\mathbf{X}+\boldsymbol{\Theta})^{T}]=-2(\mathbf{X}-\boldsymbol{\Theta})$

### v4

@@ III-C Hybrid Approach
TABLE I: Averaged simulation errors $\bar{e}_{x}$ and $\bar{e}_{y}$ for implementation of Gleser and Watson, Alg. 1, Alg. 2, and Alg. 3 with increasing standard deviation $\sigma$.
Algorithm: Standard deviation $\sigma$: 1 3 5 7 9 11 13 15
Gleser and Watson $\bar{e}_{x}$ $\mathbf{1.0975}$ $\mathbf{3.2894}$ $\mathbf{5.4887}$ $\mathbf{7.6816}$ $\mathbf{9.8777}$ $\mathbf{12.0760}$ $\mathbf{14.2788}$ $\mathbf{16.4682}$
Gleser and Watson $\bar{e}_{y}$ $0.6479$ $2.0236$ $3.5964$ $5.4476$ $7.5847$ $10.0661$ $12.8498$ $15.9657$
Algorithm 3 Hybrid solution: Implementation of a combination of the simple MLE and Gleser and Watson
@@
1:  $[\mathbf{V},\mathbf{D}]=\mathrm{eigs}(\mathbf{X}^{T}\mathbf{X}+\mathbf{Y}^{T}\mathbf{Y})$
2:  $\boldsymbol{\Theta}^{T}=\mathbf{V}\mathbf{V}^{T}\mathbf{X}^{T}$
3:  $\mathbf{B}=\mathbf{Y}\mathbf{X}^{T}(\mathbf{X}\mathbf{X}^{T})^{-1}$
4:  $\tilde{\mathbf{A}}=\mathbf{B}(1:p-1,1:p-1)$
5:  $\tilde{\mathbf{b}}=\mathbf{b}(1:p-1,p)$
6:  $\boldsymbol{\Theta}_{E}=\boldsymbol{\Theta}(1:p-1,:)$
IV Results
We evaluate the presented algorithms and investigate the difference between the implementation of Gleser and Watson and our denoised augmented improvement using Monte Carlo simulations. To test the applicability in practice, we provide the algorithms with data from an application for electronic noses.
IV-A Monte Carlo Simulations
We developed Monte Carlo simulations to evaluate the algorithms. In $l=1000$ simulation runs $n=1000$ data vectors $\boldsymbol{\theta}_{i}$ were generated with a data generation tool from [24]. We add a noise vector, where each element is a normally distributed random number scaled by the square root of the desired variance $\sigma^{2}$ using standard deviations $\sigma$ from $1$ to $15$. The data vectors have a dimension of $q=2$. We used a fixed transformation given by:
where $\tilde{\boldsymbol{\theta}}=[\tilde{\boldsymbol{\theta}}_{1},...,\tilde{\boldsymbol{\theta}}_{n}]$.
The solution of Gleser and Watson also estimates $\boldsymbol{\Theta}$, therefore, we also calculate the error defined in (16):
Finally, we calculate the arithmetic mean values of the errors $e_{x}$ and $e_{y}$ over the $l=1000$ simulation runs.
The arithmetic mean values are denoted with $\bar{e}_{x}$ and $\bar{e}_{y}$. Fig. 6 shows the mean estimation errors $\bar{e}_{x}$ and $\bar{e}_{y}$ of the algorithms. Since Algorithm 2 only estimates the parameters of the transformation between the noisy measured data from system 1 to system 2, it performs worse than Gleser and Watson taking the origins into account. In this case, the error $\bar{e}_{x}$ is only the noise of system 1. The hybrid version in Alg. 3 benefits from Gleser and Watson and delivers similar quality results for the given transformation matrix.
Analyzing error $\bar{e}_{y}$ in Table I shows the improvement of the modification. Alg. 1 improves the results of Gleser and Watson
IV-B Eigenvector Denoising by Augmentation
Fig. 5 shows the results of the Monte Carlo simulation for the eigenvector denoising extension of Alg. 1 in comparison with the Gleser and Watson standard implementation. One can see the improvement in the estimation of $\bar{e}_{y}$ while error $\bar{e}_{x}$ remains similar.
TABLE II: Normalized transformation error $\bar{e}_{y}$ linearly scaled ($min=0.00072817$ resulting 0, $max=0.0028$ resulting 1) using real data. Averaged comparison for each sensor dataset with the transformed data from all sensors for the implementation of Gleser and Watson, Alg. 1, Alg. 2, Alg. 3, and the feature-wise normalized dataset.
@@ IV-B Eigenvector Denoising by Augmentation
Table I gives an overview of the mean errors $\bar{e}_{x}$ and $\bar{e}_{y}$ for the implementation of Gleser and Watson, Alg. 1, Alg. 2, and Alg. 3 for the selected values of the standard deviation $\sigma$.
IV-C Case Study: 8 Sensor Board Scenario
For the collection of real test data we measured effects of humidity, temperature and disturbances in the air. In addition, we heat the samples during data recording and clean the sensors at regular intervals.
Data Preparation: The Bosch BME688 development kit board, shown in Fig. 3, is equipped with eight identical BME688 metal oxide sensors that are used for data recording. For the experiment, we prepared two samples [27], water and apple juice, each with a volume of 100 ml. We present only the apple juice investigations for clarity. The results with water are similar and are available in the gitlab repository. Before each measurement of the sample, the sensor is cleaned in air. The measurement period per glass is 20 minutes. The temperature of the sample glass is increased by 1 °C every 2 minutes during the measurement. The total temperature increases from 25 °C at the beginning of the measurement, to 35 °C at the end of the measurement period. The Bosch BME688 is operated with a continuously repeating temperature cycle of 10 heating temperature steps, also called parallel mode. In this example, the temperature cycle is set to 5 steps at 200 °C and 5 steps at a sensor temperature of 400 °C, each at an interval of 1260 milliseconds. For later evaluation of the measurement data, the arithmetic mean is calculated from the 5 raw values at 200 °C and 400 °C. As already mentioned in section I, Fig. 2 shows the results of the estimated transformation for an normalized example data set in the use-case of an electronic nose measuring apple juice. The black dots show the position of the data points in system $1$, blue dots represent the measured data in system 2 and the red dots are the estimation of system 2 based on the calculated transformation of Proposition 1. It can be seen that Proposition 1 allows a good estimation of the transformation in this case.
Data Evaluation: We calculated the error ${e}_{y}$ for the $64$ combinations of sensors denoting $j$ as the source sensor and $k$ as the target sensor, $j,k\in\{1,...,8\}$ by (17)
where the $\tilde{\mathbf{A}}_{j,k}$ and $\tilde{\mathbf{b}}_{j,k}$ are representing the transformation parameters from source sensor $j$ to target sensor $k$, including cases $j=k$, using $K=8$ sensors.
@@ IV-C Case Study: 8 Sensor Board Scenario
The results of the transformation error $\bar{e}_{y}$ for the normalized data set of the multi-sensor board are shown in fig. 7. The mean error for the $n=90$ data samples is calculated by (18)
$\bar{e}_{y}(j)=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{n}\sum_{i=1}^{n}\lVert(\tilde{\mathbf{A}}_{j,k}{\mathbf{x}_{j}(i)}+\tilde{\mathbf{b}}_{j,k})-\mathbf{y}_{k}(i)\rVert_{2}.$ (18)
Since we can use the measured values $\mathbf{y}_{k}$, we do not need to calculate them as in eq. 15 by $(\mathbf{A}\boldsymbol{\theta}_{i}+\mathbf{b})$. Table II shows the normalized mean error $\bar{e}_{y}$ for each source sensor, with the minimum value $0.00072817$ resulting 0, and maximum value $0.0028$ resulting 1. Since we focus on calculating the transformation from $\mathbf{X}$ to $\mathbf{Y}$, and do not try to estimate the origins of the data, the performance of Gleser and Watson is slightly less accurate than the least squares, and the hybrid version delivers similar results as Proposition 1. Our proposed Alg. 2, with reduced computational steps and therefore lower energy consumption, gives the best results except source sensor 6. However, this could also be due to inaccuracies in the measurement procedure and needs further investigation. A simple normalization of the data in our experiment leads to at least $50~{}\%$ worse results than our proposed method.
@@ V Discussion and Outlook
In the article, we have shown that the procedures of estimating the 2-dimensional AT can be used for an expert-assisted learning between two systems. The simple direct estimation based on the stationary points of the parameters of the transformation can be combined with the solution for the estimation of the data from Gleser and Watson. This new hybrid solution provides the best estimates for both, simulations and real data. For the presented real data evaluation, we achieve best results with the simple Prop. 1, but we are looking forward to investigate whether we can use the approach of Gleser and Watson to improve the measurement values of our low-cost sensors in an extended setup. For the real data example, our simple least squares approach already achieves unique solution due to the non singularity in $\boldsymbol{\Theta}\boldsymbol{\Theta}^{T}$, and the best accuracy with low computation, which is very promising for further developments in distributed learning. The results still need to be generalized for the m-dimensional case.
Finally, in further work we will explore if only a part of the measured data is sufficient to give a good estimate of the transformation. The evaluation with multiple transformation can also be done in future work. Based on this, iterative methods can be developed in the future to estimate the transformation between different systems. Fig. 8 illustrates the concept of our future work to implement distributed learning based on the creation of similar models using the expert-supported transformation estimation (see the figure in the abstract) between the two data spaces. As to our knowledge, there is no similar approach, although it seems to be a simple way to match distributed systems and their events or anomalies. We plan to implement these concepts as the basis for a more efficient method of knowledge sharing between distributed nodes. We also plan to investigate in more detail why denoising improves the results. In our example, the fixed transformation we have chosen is favorable for Gleser and Watson However, when considering various alternatives for $\mathbf{A}$, Gleser and Watson produces notably inferior results. We encourage readers to experiment with this option and assess its performance in different $\mathbf{A}$ scenarios using a code provided on GitLab at https://gitlab.rlp.net/vski-engineering-group/paper/likelihood-based-sensor-calibration-for-expert-supported-distributed-learning-algorithms-in-iot-systems-estimating-affine-transformations/matlabcodeanddata.
@@ VI Conclusion
We propose an improved solution of Gleser and Watson by augmentation of the matrix, denoising using the eigenvectors and the extension to an AT. We have shown that both solutions are connected by an eigenvalue decomposition. We further showed that a combination of both solutions provides better estimates for computing the parameters of the AT by both, simulation, and a real experiment. Finally we presented a novel concept for distributed learning using transformation with the support of expert knowledge.
All 3 algorithms offer a good approach to estimate the AT between similar systems. Further investigations should include special methods to fairly compare Gleser and Watson with our hybrid proposition. The development of a simple but powerful expert-based distributed learning algorithm using ATs appears promising but requires further research in multidimensionality, and an iterative approach. Investigating more powerful transformation for stronger IoT nodes should make sense as well.
@@
$\nabla_{\boldsymbol{\Theta}}\mathrm{tr}[(\mathbf{X}-\boldsymbol{\Theta})(\mathbf{X}-\boldsymbol{\Theta})^{T}]=-2(\mathbf{X}-\boldsymbol{\Theta})$